OpenStudy (anonymous):

1. If a snail is traveling at a rate of 1 inch per minute, how many miles per hour is the snail traveling? Give your response using scientific notation

OpenStudy (austinl):

\[(1inch/1\min)\times(1foot/12 inches)\times(1mile/5280 feet)\times(60 minutes/1 hour)=9.47\times10^-4 mph\]

OpenStudy (austinl):

It got cut off... \[9.47\times10^{-4} mph\]